Startup Night 1 - 20140830 UT

Observers: JHill (LBTO), OKuhn, BRothberg (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support: JChristou
Telescope Support: KNewton
Software Support: KSummers, DSummers
Instrument Support: JPower (LBTO)
Engineering Support: DAshby

Summary

LBC

1) Red, but not Blue has problems turning the power on. The UI won't turn on Red housekeeping. It seems even worse than it was earlier in the night. I experience a similar problem with the primitive commands: powstatus, etc. They work for blue, but fail frequently for red. See the core dumps at 08:17. This may have been a funky Dataprobe.

1a) Red still has complaints of a shutter hardware error that we were blaming on power problems, but maybe not.

2) Blue can do everything but guide. The lbc software crashes just after it sends the first non-zero guide correction to the telescope. (repeated 3 times to be sure it is always at that place). If I uncheck the "Guiding" box, then it takes tech chip images but doesn't send corrections to the telescope, and everything is fine.

3) The various problems you saw with the UI getting hung up and having delayed messages.

MODS1

MODS1 seems to be working OK after the daytime repairs. Of course, SX AdSec is still RIP'd, so we never got decent images.

Preparations

Worked all summer........

Details

SX AdSec will not turn on. Mike L and Juan Carlos are coming up on Saturday.

James R and Jenny are working on repairing LBC cables, so we will open late.

Polycom complains about not talking to Gatekeeper (from here and from N525, but both units can dial N505 ??)

03:34 James R departs the mountain.

Sky is magnificently clear, and will be dark as soon as the crescent moon sets. Wind is light from SW. T=12 degC D=2.5 degC.

04:10 Steve is trying to open,

04:27 Left Shutter Door is Open, Problems with balls on Latch Pin 0708. See Issue 5279. Then the right upper shutter door drive SD1002 needed a reset.

04:30 LBC-Red Shutter is not responding correctly, although Kellee can talk to it on minicom.

04:36 Right Shutter Door is Open. Moving to Zenith to fill MODS1. Moving LBC swing arms in.

LBC Wake-up

04:48 LBC-Blue bias successful. Taking 10 more.

04:56 Turn LBC off, so we can turn on Red again. Kellee reset the Red shutter controller through minicom.

04:58 Turn LBC on with both sides.

05:12 Reboot the new CMU because problems with the UI not displaying.

05:16 Steve completes a preset from IRAF ACT0289.

05:20 Continuing problems with LBC-Red power control. So back to using LBC-Blue alone.

05:24 Preset from LBC ACT0289.

05:30 DOFPIA, /BLUEONLY

05:39 180 sec ACT0289

05:40 LBC program crashed and restarted after the first guide correction.

2014/08/30 05:40:54.223523 N B ROTATOR           achieved PositionAngle 0.0501 deg, PA diff -0.0012, RotatorAngle -44.2414
2014/08/30 05:40:54.376494 N B ROTATOR           GOYA query:>CTRL_A< answer:>D32X65|<
2014/08/30 05:40:54.513206 N B ROTATOR           GOYA query:>CTRL_E< answer:>X0.00439 |<
2014/08/30 05:40:54.649876 N B ROTATOR           GOYA query:>CTRL_I< answer:>X0.62998 F0.62998|<
2014/08/30 05:40:54.786617 N B ROTATOR           GOYA query:>CTRL_W< answer:>X0.58008 |<
2014/08/30 05:40:55.237363 N B ROTATOR           achieved PositionAngle 0.0519 deg, PA diff -0.0029, RotatorAngle -44.2326
2014/08/30 05:40:55.242306 N B TRACKERS          TCS IIF_OffsetGuiding completed


2014/08/30 05:41:05.287052 < Log Service: session started >
2014/08/30 05:41:05.287211 < Log Service: RPC server started >
2014/08/30 05:41:05.288096 N -                   Starting LBC software version 2.0 - Aug 29 2014
2014/08/30 05:41:05.288151 N -                   clear ProfiBus control handles
2014/08/30 05:41:05.292993 N -                   local data initialize ...
2014/08/30 05:41:05.293084 N -                   local data initialize completed
2014/08/30 05:41:05.293109 N -                   accessing /lbccontrol/current/conf/lbc.conf configuration file ...
2014/08/30 05:41:05.293366 S -                   read ENABLE parameter >LBC< with value TRUE
2014/08/30 05:41:05.293400 S -                   read ENABLE parameter >LBC.UseTelescope< with value TRUE

05:52 Turning LBC-Blue on again.

05:57 Play 180x OB on ACT0289 and it crashed again after the first guide correction.

2014/08/30 05:57:37.596276 N B ROTATOR           GOYA query:>CTRL_A< answer:>D32X97|<
2014/08/30 05:57:37.732977 N B ROTATOR           GOYA query:>CTRL_E< answer:>X0.00220 |<
2014/08/30 05:57:37.869832 N B ROTATOR           GOYA query:>CTRL_I< answer:>X0.98104 F0.98104|<
2014/08/30 05:57:38.006413 N B ROTATOR           GOYA query:>CTRL_W< answer:>X0.73389 |<
2014/08/30 05:57:38.517745 N B ROTATOR           achieved PositionAngle 0.0711 deg, PA diff -0.0052, RotatorAngle -30.8380
2014/08/30 05:57:38.650994 N B ROTATOR           GOYA query:>CTRL_A< answer:>D32X97|<
2014/08/30 05:57:38.787618 N B ROTATOR           GOYA query:>CTRL_E< answer:>X0.00220 |<
2014/08/30 05:57:38.907199 N B TRACKERS          TCS IIF_OffsetGuiding completed


2014/08/30 05:57:48.937403 < Log Service: session started >
2014/08/30 05:57:48.937573 < Log Service: RPC server started >
2014/08/30 05:57:48.938440 N -                   Starting LBC software version 2.0 - Aug 29 2014
2014/08/30 05:57:48.938494 N -                   clear ProfiBus control handles
2014/08/30 05:57:48.943457 N -                   local data initialize ...
2014/08/30 05:57:48.943540 N -                   local data initialize completed
2014/08/30 05:57:48.943556 N -                   accessing /lbccontrol/current/conf/lbc.conf configuration file ...
2014/08/30 05:57:48.943808 S -                   read ENABLE parameter >LBC< with value TRUE
2014/08/30 05:57:48.943844 S -                   read ENABLE parameter >LBC.UseTelescope< with value TRUE
2014/08/30 05:57:48.943866 S -                   read ENABLE parameter >LBC.BlueChannel< with value TRUE
2014/08/30 05:57:48.943895 S -                   read STRING parameter >LBC.BlueChannel.FileNameLeader< with value >lbc

06:05 DOFPIA, /BLUEONLY

06:09 RB_rV_superfoc7

06:14 RB_rV_superfoc7_pa180

06:22 DOFPIA, /BLUEONLY

06:26 RB_rV_superfoc7

06:34 RB_rV_superfoc7_pa180

06:42 OSU_monitor_N6946focus

some thin cirrus clouds are drifting in from the north.

06:48 OSUmonitorN6946 (with guiding turned off [unchecked on the UI], still taking guide images) - This did not crash the lbc program. The 300 sec unguided images are round.

07:24 replay the same OB with guiding checked. Crashed after the first guide guide expose.

MODS1

07:27 Move to zenith and switch the left side to MODS1.

07:42 Steve sends a MODS1 ACQUIRE preset from IRAF LBTtools.

Trying to collimate MODS1 while the shell is ripped doesn't do all that well.

08:08 John accidentally powers off the new CMU. (Don't type "poweroff", but rather ./powoff.)

08:15 Test of M31 guidestar. It looks OK against the background of the galaxy.

Core dumps near 8:17 are from ./powstatus redhk:status (Red also fails for powstatus, but blue seems ok.)

08:28 Moving to Feige 110. Olga and Barry ran a spectrophotometric acquisition & obtained data in dual/red/blue grating/prism, but the telescope was not able to collimate (secondary could not be powered on) and the star overfilled the 5" slit, so these data are not to be used for any flux calibrations (the image title indicates these are test data).
  • acquisition OK
    • acqMODS feige110
    • modsAlign -r
    • uncollimated star fills 5" slit
  • run through all six modes, 1 spectrum per each.
    • star appears double in some prism spectra due to runaway "collimation"

LBC-Blue

09:02 Switching back to LBC-Blue to take some more photometric monitoring fields.

09:09 warm reboot of cmu.

Nothing worked until we realized that stuck Dataprobe 123 (RDprobe1) was causing LBC software to crash every 3 minutes.

While John was working to get LBC running, obtained two sets of data to remeasure/verify collimator TTF zeropoints.

10:30 Steve moved telescope to 60 deg elevation and LDG rotator at 207 deg.
  • execMODS imcsTTFs.pro
    • ran twice, image numbers b 22-27, 28-33 and r 41-46, 47-52.
  • examination against examples show little change in zeropoints, as expected.
  • Values are not yet in place on mods1 computer --- TBD in the afternoon.

11:30 Now back to LBC UI working, but LBC-Red still reporting a shutter hardware error.

DOFPIA

11:35 N672focus

DOFPIA

11:39 OSUmonitorN672.ob but I forgot to uncheck guiding, so it crashed.

11:48 Close for impending dawn.

-- %USERSIG{JohnHill - 2014-08-30}%
Topic revision: r5 - 31 Aug 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback